How are you going to prepare 1.5 L of 0.1N KOH solution? Molar mass of KOH = 56 g/mol.
1
Expert's answer
2020-09-24T11:21:20-0400

 1.5 L of 0.1N KOH solution Have moles of KOH = 1.5 * 0.1 = 0.15 mol .


Mass of KOH = 0.15 * 56 = 8.4 g . Answer .